# Towards Learning Abductive Reasoning using VSA Distributed Representations

### v1

@@ 1 Introduction
An alternative and promising direction is neuro-symbolic AI. Neuro-symbolic approaches combine sub-symbolic perception with various forms of symbolic reasoning, resulting in cutting-edge performance across a spectrum of domains, including visual [23, 24, 25, 26], natural language [27], causal [28], mathematical [29], and analogical [30, 31, 32, 33, 34] reasoning tasks. In the context of RPM, recent neuro-symbolic architectures focused on abductive reasoning [33, 34]. Abductive reasoning allows to selectively infer propositions based on prior knowledge represented in a symbolic form to explain the perceptual observations in the best possible way [35]. The appeal of the abductive approach lies in its accommodation of perceptual uncertainties within symbolic reasoning. Abductive reasoning can be implemented in systems that leverage distributed vector-symbolic architectures (VSAs) [36, 37, 38] representations and operators, such as the Neuro-Vector Symbolic Architecture (NVSA) [34]. However, these neuro-symbolic architectures [33, 34] necessitate complete knowledge over the application domain (which might not be available) to program the right inductive bias into the model.
@@
Learn-VRF [39] overcomes this limitation by introducing a probabilistic abduction reasoning approach that learns a subset of the rules underlying RPM from data. Learn-VRF transparently operates in the rule space, learning them through a soft-assignment of VSA attribute representations to a fixed rule template. During inference, it generates the answer panel by executing all the learned rules and applying a soft-selection mechanism to their outputs. Nevertheless, Learn-VRF comes with several limitations, including a sub-optimal selection mechanism, poor performance on the RPM constellations involving multiple objects, and a constraint on the expressiveness of the RPM rules it can learn.
@@
To make progress towards learning-to-reason, we propose the Abductive Rule Learner with Context-awareness (ARLC) to tackle the main limitations of Learn-VRF [39]. We advance a novel context-augmented formulation of the optimization problem and a more expressive rule template, which allow sharing rules with the same parameters in both execution and selection steps and offer better interpretability. An overview of ARLC is depicted in Figure 1. ARLC features programmability and can further learn from data on top of programmed knowledge. We evaluate ARLC on in-distribution (ID) and out-of-distribution (OOD) tests of the I-RAVEN dataset, and demonstrate that ARLC significantly outperforms neuro-symbolic and connectionist baselines, including large language models. Further, the number of trainable parameters is reduced by two orders of magnitude compared to Learn-VRF. We experimentally validate the programmability of ARLC by encoding domain knowledge, and discover that post-programming training, contrary to other studies [40], does not compromise the validity of the solution, but rather improves it. Finally, training the model on a single constellation and evaluating it on all the others, we show that, unlike previous baselines [39, 34], the learned rules can seamlessly transferred across constellations of the I-RAVEN dataset.
@@ 2.1 Vector Symbolic Architectures
Bundling is a similarity-preserving operation that creates a superposition of the operands, that is, the resulting vector will have a high similarity with the two operands. Binding, on the other hand, is an operation that allows to bind a vector (value) to another vector (key) and does not preserve similarities; it usually allows an inverse operation, called unbinding. The specific realization of the bundling, binding, and vector space constitute the main difference between members of the VSA family.
@@ 2.3 Learning to Reason with Distributed Representations
In this section, we discuss how vector-symbolic architectures can be used to solve tasks which require analogical and relational reasoning such as RPM. In particular, we focus on Learn-VRF [39], a simple yet powerful approach that enables to solve RPM tests by learning the underlying relations between visual attributes in the VSA representational space.
The key observation behind this approach is that the formulation of every RPM rule in the VSA algebra is a particular instance of a general rule template, which is shared among all the rules and consists of a series of binding and unbinding operations between VSA vectors. Hence, the problem of learning RPM rules can be framed as an assignment problem between vectors representing visual attributes and terms in this general rule template. This alternative formulation allows to tackle one of the main limitations of neuro-symbolic approaches, differentiability, and therefore enables the use with data-driven learning algorithms based on gradient optimization.
Learn-VRF includes several sequential steps, ranging from the translation of visual attributes into the VSA high-dimensional space to the computation of final results, that are detailed in the following paragraphs.
@@
Inference with the learned rule set is a two-steps process: an execution step (where all the rules are applied in parallel to the input) and a selection step (where a prediction for the missing panel is generated). The application of each rule to an RPM example generates a tuple of three VSA vectors , which corresponds to the result of the rule execution on the three rows of the RPM matrix, together with a rule confidence value . The confidence value is computed as the sum of the cosine similarities between the predicted VSA vectors and their respective ground-truth vector,
During inference, the last term of the sum () is omitted, as the ground-truth for the third row is unknown.
@@
The use of the weighted combination can be understood as a soft selection mechanism between rules, and was found to be more effective compared to the hard selection mechanism provided by sampling [39].
@@ 3 Methods
In this section, we present our Abductive Rule Learner with Context-awareness (ARLC) system. An overview of ARLC is depicted in Figure 1. The framework aligns with the original Learn-VRF at a conceptual level, albeit with key adjustments that allow to improve its expressiveness and boost its downstream performance on the I-RAVEN dataset.
@@ 3.1 Learning Context-Augmented RPM Rules
To overcome this limitation, Learn-VRF instantiates and learns three different rule sets (one per row) simultaneously. During inference, the model leverages the first two sets to produce confidence values (Equation 3), which are then used to perform a soft-selection of the output panels produced by the third rule set (Equation 4). While it was empirically shown to be effective, this implementation leaves the door open to different criticalities. For instance, the model has no constraint on the functional equivalence between the three learned rule sets. This renders the interpretability of the model sensibly harder and increases the likelihood of learning spurious correlations in the rule selection mechanism. Furthermore, this formulation diminishes the model’s versatility, as its design, tailored to the RPM context, cannot seamlessly transfer to other abstract reasoning tasks without a reconfiguration of its primary components.
@@ 3.3 Training Loss and other Implementation Aspects
We follow the training recipe provided by Learn-VRF [39]. The training loss is defined as the inverse cosine similarity between the three predicted panels and their corresponding ground-truth
@@ 4.1 In-distribution (ID) Results
Contrary to all the other methods (with the exception of PrAE [33]), ARLC is exclusively trained on the 2x2 constellation, effectively reducing the number of trained parameters and training samples by 85% (). Its seamless adaptation to unseen constellations demonstrates the generality of the learned rules but prevents the model from outperforming the baselines in each single constellation.
@@ 5 Conclusions and Future Work
A potential avenue for future research involves addressing the remaining challenge posed by this dataset, specifically, the development of suitable representations to allow the learnability of arithmetic and progression rules on the position attributes. These rules are currently impossible to learn for the model and would allow it to attain perfect accuracy on I-RAVEN. Additionally, extending the evaluation of ARLC to include other reasoning benchmarks, such as ARC [3], also represents a promising direction for further investigation. In fact, while the scope of this work was mostly focus on developing and studying a prototype that could solve RPM, the proposed general formulation of the problem could transfer to other settings that require relations and analogies data-driven learning. Furthermore, as our experiments on learning of top programmed knowledge show, our framework holds potential for application in scenarios where only a partial knowledge on the dynamics is available, and it is necessary to discover and build new knowledge on top of it.

### v3

@@ 1 Introduction
An alternative and promising direction is neuro-symbolic AI. Neuro-symbolic approaches combine sub-symbolic perception with various forms of symbolic reasoning, resulting in cutting-edge performance across a spectrum of domains, including visual [23, 24, 25, 26], natural language [27], causal [28], mathematical [29], and analogical [30, 31, 32, 33, 34] reasoning tasks. In the context of RPM, recent neuro-symbolic architectures focused on abductive reasoning [33, 34]. Abductive reasoning allows to selectively infer propositions based on prior knowledge represented in a symbolic form to explain the perceptual observations in the best possible way [35]. The appeal of the abductive approach lies in its accommodation of perceptual uncertainties within symbolic reasoning. Abductive reasoning can be implemented in systems that leverage distributed vector-symbolic architectures (VSAs) [36, 37, 38] representations and operators, such as the Neuro-Vector Symbolic Architecture (NVSA) [34]. However, these neuro-symbolic architectures [33, 34] necessitate complete knowledge of the application domain (which might not be available) to program the right inductive bias into the model.
@@
Learn-VRF [39] overcomes this limitation by introducing a probabilistic abduction reasoning approach that learns a subset of the rules underlying RPM from data. Learn-VRF transparently operates in the rule space, learning them through a soft assignment of VSA attribute representations to a fixed rule template. During inference, it generates the answer panel by executing all the learned rules and applying a soft-selection mechanism to their outputs. Nevertheless, Learn-VRF comes with several limitations, including a sub-optimal selection mechanism, poor performance on the RPM constellations involving multiple objects, and a constraint on the expressiveness of the RPM rules it can learn.
@@
To make progress towards learning-to-reason, we propose the Abductive Rule Learner with Context-awareness (ARLC) to tackle the main limitations of Learn-VRF [39]. We advance a novel context-augmented formulation of the optimization problem and a more expressive rule template, which allows sharing rules with the same parameters in both execution and selection steps and offers better interpretability. An overview of ARLC is depicted in Figure 1. ARLC features programmability and can further learn from data on top of programmed knowledge. We evaluate ARLC on in-distribution (ID) and out-of-distribution (OOD) tests of the I-RAVEN dataset and demonstrate that ARLC significantly outperforms neuro-symbolic and connectionist baselines, including large language models. Further, the number of trainable parameters is reduced by two orders of magnitude compared to Learn-VRF. We experimentally validate the programmability of ARLC by encoding domain knowledge, and discover that post-programming training, contrary to other studies [40], does not compromise the validity of the solution, but rather improves it. Finally, training the model on a single constellation and evaluating it on all the others, we show that, unlike previous baselines [39, 34], the learned rules can seamlessly be transferred across constellations of the I-RAVEN dataset.
@@ 2.1 Vector Symbolic Architectures
Bundling is a similarity-preserving operation that creates a superposition of the operands, that is, the resulting vector will have a high similarity with the two operands. Binding, on the other hand, is an operation that allows to bind a vector (value) to another vector (key) and does not preserve similarities; it usually allows an inverse operation, called unbinding. The specific realization of the bundling, binding, and vector space constitutes the main difference between members of the VSA family.
@@ 2.3 Learning to Reason with Distributed Representations
In this section, we discuss how vector-symbolic architectures can be used to solve tasks that require analogical and relational reasoning such as RPM. In particular, we focus on Learn-VRF [39], a simple yet powerful approach that enables solving RPM tests by learning the underlying relations between visual attributes in the VSA representational space.
The key observation behind this approach is that the formulation of every RPM rule in the VSA algebra is a particular instance of a general rule template, which is shared among all the rules and consists of a series of binding and unbinding operations between VSA vectors. Hence, the problem of learning RPM rules can be framed as an assignment problem between vectors representing visual attributes and terms in this general rule template. This alternative formulation allows to tackle one of the main limitations of neuro-symbolic approaches, differentiability, and therefore enables the use of data-driven learning algorithms based on gradient optimization.
Learn-VRF includes several sequential steps, ranging from the translation of visual attributes into the VSA high-dimensional space to the computation of final results, which are detailed in the following paragraphs.
@@
Inference with the learned rule set is a two-step process: an execution step (where all the rules are applied in parallel to the input) and a selection step (where a prediction for the missing panel is generated). The application of each rule to an RPM example generates a tuple of three VSA vectors , which corresponds to the result of the rule execution on the three rows of the RPM matrix, together with a rule confidence value . The confidence value is computed as the sum of the cosine similarities between the predicted VSA vectors and their respective ground-truth vector,
During inference, the last term of the sum () is omitted, as the ground truth for the third row is unknown.
@@
The use of the weighted combination can be understood as a soft selection mechanism between rules and was found to be more effective compared to the hard selection mechanism provided by sampling [39].
@@ 3 Methods
In this section, we present our Abductive Rule Learner with Context-awareness (ARLC) system. An overview of ARLC is depicted in Figure 1. The framework aligns with the original Learn-VRF at a conceptual level, albeit with key adjustments that improve its expressiveness and boost its downstream performance on the I-RAVEN dataset.
@@ 3.1 Learning Context-Augmented RPM Rules
To overcome this limitation, Learn-VRF instantiates and learns three different rule sets (one per row) simultaneously. During inference, the model leverages the first two sets to produce confidence values (Equation 3), which are then used to perform a soft selection of the output panels produced by the third rule set (Equation 4). While it was empirically shown to be effective, this implementation leaves the door open to different criticalities. For instance, the model has no constraint on the functional equivalence between the three learned rule sets. This renders the interpretability of the model sensibly harder and increases the likelihood of learning spurious correlations in the rule selection mechanism. Furthermore, this formulation diminishes the model’s versatility, as its design, tailored to the RPM context, cannot seamlessly transfer to other abstract reasoning tasks without a reconfiguration of its primary components.
@@ 3.3 Training Loss and other Implementation Aspects
We follow the training recipe provided by Learn-VRF [39]. The training loss is defined as the inverse cosine similarity between the three predicted panels and their corresponding ground truth
@@ 4.1 In-distribution (ID) Results
Contrary to all the other methods (except for PrAE [33]), ARLC is exclusively trained on the 2x2 constellation, effectively reducing the number of trained parameters and training samples by 85% (). Its seamless adaptation to unseen constellations demonstrates the generality of the learned rules but prevents the model from outperforming the baselines in each single constellation.
@@ 5 Conclusions and Future Work
A potential avenue for future research involves addressing the remaining challenge posed by this dataset, specifically, the development of suitable representations to allow the learnability of arithmetic and progression rules on the position attributes. These rules are currently impossible to learn for the model and would allow it to attain perfect accuracy on I-RAVEN. Additionally, extending the evaluation of ARLC to include other reasoning benchmarks, such as ARC [3], also represents a promising direction for further investigation. While the scope of this work was mostly focused on developing and studying a prototype that could solve RPM, the proposed general formulation of the problem could transfer to other settings that require learning relations and analogies from data. Furthermore, as our experiments on learning of top programmed knowledge show, our framework holds potential for application in scenarios where only a partial knowledge of the dynamics is available, and it is necessary to discover and build new knowledge on top of it.